OpenStudy (santashelper):

Your friend comes up with a design for the perfect garden and says to create it you need an area of 35 square feet. if this is your ideal garden, what would be the measurements of your width and length

OpenStudy (anonymous):

This has a many multitude of possibilities

OpenStudy (santashelper):

you solve with a quadratic equation if that helps!

OpenStudy (anonymous):

First, the equation of an area: A= w*l, where "w" is width and "l" is length. So we have 35 = w*l. In other words, think of two numbers that when multiplied together equals 35!

OpenStudy (santashelper):

5 and 7?

OpenStudy (anonymous):

That'll work. That way, you get an area of 35. Specifically it should be 5ft and 7ft if your teacher is really into that.

OpenStudy (santashelper):

how do i show it on a Quad equation?

OpenStudy (santashelper):

also, the second part of the question is….during the winter months, MR Dugan assigns another set of students to design a rectangular tent to cover your garden. the volume of their tent was represented by 2x^2-5x^2+3x. the garden still had the original length and width from the previous example, what was the measure of height in terms of x

OpenStudy (santashelper):

^ second part involves factoring

OpenStudy (anonymous):

Well, if you need to use a quadratic equation, it changes things. One, it ensures the garden has to be a square. Because it's a square, all the sides are equal, in other words, w = l. So, 5ft and 7ft are no longer acceptable answers.

OpenStudy (santashelper):

wow youre awesome… i didnt even realize… how do i go about it now

OpenStudy (anonymous):

So, you now have the w^2 (or l^2. I'm using w). so w^2 = A. The quadratic equation is ax^2 + bx + c = 0. 'x' typically represents the unknown, which in this case will be 'w' (keep in mind that a ≠ A). So aw^2 + bw + c = 0. If the original area equation is w^2 = 35, we need to alter it to match the quadratic formula. First, we subtract 35 from both sides. This gets us w^2 - 35 = 0 So then c has to be -35. So what to a and b have to be to match that equation?

OpenStudy (anonymous):

Also, keep in mind that since this is something in the real world, w CANNOT be negative. Otherwise your ideal garden wouldn't exist.

OpenStudy (santashelper):

so do you find the square root of 35 for the answer

OpenStudy (anonymous):

Seems so. Using the quadratic equation seems really roundabout though.

OpenStudy (santashelper):

okk i will just do that… do you know how to do part two?? ….during the winter months, MR Dugan assigns another set of students to design a rectangular tent to cover your garden. the volume of their tent was represented by 2x^2-5x^2+3x. the garden still had the original length and width from the previous example, what was the measure of height in terms of x

OpenStudy (santashelper):

using factoring

OpenStudy (anonymous):

Alright so we have the volume of a rectangle being (w*l*h). The equation given for the prism is 2x^2 - 5x^2 +3x. First things first, simplify it a little to -3x^2 + 3x. So the two volumes have to equal each other, right? so -3x^2 + 3x = (w*l*h) As we discovered in question 1, w (and by default l) equal sqrt(35). So it should be -3x^2 + 3x = (35 * h), making it 3x^2 + 3x = 35h. In quadratic terms, it's -3x^2 + 3x -35h = 0

OpenStudy (anonymous):

You can't factor with h in the equation though, so you have to move it back. All factoring really does here is simplify the equation. So what do -3x*x and 3*x have in common? Take whatever they have in common out and see what remains. Don't forget to make sure h is alone at the end.